OpenStudy (anonymous):

Find all t that satisfy cos(t)=-1/2 for t belongs to [0,3pi]

OpenStudy (aum):

In which quadrants is cosine negative?

OpenStudy (anonymous):

Is it II and IV?

OpenStudy (aum):

It is II and III

OpenStudy (aum):

cos(pi/3) = 1/2. So the reference angle is pi/3. Since cosine is negative in quadrant II and III, the angles are: pi - pi/3, pi + pi/3, for angles in [0,2pi]. Simplify the angles first and then we will get to [0,3pi].

OpenStudy (anonymous):

2pi/3 and 4pi/3?

OpenStudy (aum):

correct. Now 3pi = 2pi + pi. So we need one more angle that falls in the second quadrant. Out of 2pi/3 and 4pi/3, 2pi/3 is in the second quadrant to which if we add a 2pi it will make one full rotation and come back to quadrant II. 2pi/3 + 2pi = 2pi(1/3 + 1) = 2pi * 4/3 = 8pi/3. Therefore, in the interval [0, 3pi], the angles are 2pi/3, 4pi/3, 8pi/3

OpenStudy (anonymous):

I see! Thank you so much!!!

OpenStudy (aum):

You are welcome. The above answers in radians is the correct way to answer it. But for ease of understanding, the angles in degrees are: 120, 240, 480.